An investment of $10,000 can be made in a project that will produce five (5) uniform annual revenues and a market value (salvage) value of $2,000. The company is willing to accept any project that will earn 10% per year or more, on all the invested capital. Compute the annuity that will make this investment worthwhile.

$2,000

$1,600

$2,310

$2,638
Answer #1

Let annuity value be A

AW=-10000(A/P,10%,5)+A+2000(A/F,10%,5)

For the investment to be worthwhile

Annuity = 10000(A/P,10%,5)-2000(A/F,10%,5)

=10000(0.263797)-2000(0.163797)

=2637.97-327.594

=2310.37

Option(C)